# A Weak Supervision Learning Approach Towards an Equitable Parking Lot Occupancy Estimation

### v1

@@ Introduction
Figure 1: (Top row) A PlanetScope imagery showing the parking situation of a large parking lot observed on a typical Saturday (left), a typical Sunday (right). (Bottom row): A PlanetScope image showing the parking occupancy situation on a small parking lot observed on a Saturday (left), and a Sunday (right)
@@
Figure 2: Data acquisition pipeline from POI identification on OSM to acquiring PlanetScope images
@@
[…]
@@
Figure 3: Given a pair of images, Encoder 1 and Encoder 2 are used to extract features of a lower dimension, yielding Feature 1 and Feature 2 for image 1 and image 2 respectively. The difference between the feature vectors is then projected to a single output with a Multi-Layer Perceptron (MLP)
@@
Table 1: AUC score of pairwise comparison for different parking lot sizes
In our final experiment, we seek to validate our approach in the context of the on-going conflict between the Sudanese Armed Forces and the Rapid Support Forces that broke out in April 2023. As reported by Guo et al. (2024), there were fewer cars in Sudan during the post_war period (April 14–21, 2023) compared to the pre_war period (April 1–7, 2023), as reflected by a decrease in nitrogen dioxide ($\text{NO}_{2}$) concentrations, which are primarily emitted by vehicles and power plants. To assess whether our method reflects this observation, we analyzed PlanetScope images of the Jackson Bus Terminal in Khartoum. Specifically, we select days of the week in both the pre_war and post_war periods for which images were available, then apply our pairwise comparison approach to all possible image pairs (excluding pairs with the same date). As shown in Figure 4, images captured during the war period consistently received lower rankings than their pre_war counterparts, aligning with the observed drop in $\text{NO}_{2}$ concentrations.

### v2

# A Weak Supervision Learning Approach Towards an Equitable Mobility Estimation

@@ Abstract
 [^1] To appear in the proceedings of the ICWSM’25 Workshop on Data for the Wellbeing of Most Vulnerable (DWMV). Please cite accordingly.
@@ Introduction
Figure 1: (Top row) A PlanetScope imagery showing the parking situation of a large parking lot observed on a typical Saturday (left), a typical Sunday (right). (Bottom row): A PlanetScope image showing the parking occupancy situation on a small parking lot observed on a Saturday (left), and a Sunday (right).
@@
Figure 2: Data acquisition pipeline from POI identification on OSM to acquiring PlanetScope images.
@@
After these preprocessing steps, we retain $683$ parking lots which include $355$ large parking lots ($\geq 10,000$ sqm) , $142$ medium parking lots (between $5,000$ sqm and $10,000$ sqm) and $186$ small parking lots ($\leq 5,000$ sqm).
@@
Figure 3: Given a pair of images, Encoder 1 and Encoder 2 are used to extract features of a lower dimension, yielding Feature 1 and Feature 2 for image 1 and image 2 respectively. The difference between the feature vectors is then projected to a single output with a Multi-Layer Perceptron (MLP).
@@
Table 1: AUC score of pairwise comparison for different parking lot sizes.
In our final experiment, we seek to validate our approach in the context of the on-going conflict between the Sudanese Armed Forces and the Rapid Support Forces that broke out in April 2023. As reported by Guo et al. (2024), there were fewer cars in Sudan during the post_war period (April 14–21, 2023) compared to the pre_war period (April 1–7, 2023), as reflected by a decrease in nitrogen dioxide (${NO}_{2}$) concentrations, which are primarily emitted by vehicles and power plants.
To assess whether our method reflects this observation, we analyzed PlanetScope images of the Jackson Bus Terminal in Khartoum. Specifically, we select days of the week in both the pre_war and post_war periods for which images were available, then apply our pairwise comparison approach to all possible image pairs (excluding pairs with the same date). As shown in Figure 4, images captured during the war period consistently received lower rankings than their pre_war counterparts, aligning with the observed drop in ${NO}_{2}$ concentrations.